OpenStudy (anonymous):

Use the laws of logarithms to express each side of the equation as a single logarithm. Then compare both sides of the equation to solve. Please help me. Thank you :)

OpenStudy (anonymous):

\[\log _{2}x= 2\log _{2}7 +\log _{2}5\]

OpenStudy (loser66):

your attempt first

OpenStudy (anonymous):

The way I did it, x=245. I'm not sure if i did it correctly

OpenStudy (loser66):

how to get?

OpenStudy (anonymous):

 is this the correct process?